th oduced 15.is Fe oxcess Car 3 s the actual yiald of iron in moles? uhat is ine theoretical iadt of iran in moles2 hhat is the percent ye
Answer #1

1. The molar weight of iron is 56 grams.

Actual yield of iron in moles=15.7/56 moles = 0.28 moles

2. In the above reaction 1 mole Fe203 produce 2 mole Fe. Hence, 0.225 mole Fe203 produce 0.45 mole Fe. So, the theoretical yield of Fe= 0.45 mole

3. The pecent of yield= 0.28/0.45 *100 = 62.22%